An Electron revolves in a circular orbit of radius ‘r’ with angular speed ‘ω’. Using the expression for the magnetic field due to a circular current loop, deduce an expression for the magnetic field at the center of the Electron orbit.

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
magnetic field due to a circular coil at the centre is
B = \mu0 I / 2R
I = q/t = e * (2*pi)/ w
Hence
B = \mu0 * e / Rw

Regards
Arun (askIITians forum expert)